# Optimally-Weighted Maximum Mean Discrepancy Framework for Continual Learning

### v1

@@ 4.1 Experiment Setup
Implementation Details. We employ ResNet18 as our foundational architecture, comprising five layers dedicated to feature extraction. To derive the adaptive regularization optimization term, we initialized five adaptive weights randomly, sourced from a uniform distribution. Our methodology extends the code from Refresh Learning [54], integrating our approach within their established framework. We utilized the hyperparameters specified in their implementation and conducted a grid search to ascertain the optimal value for the $\gamma$ hyperparameter, aiming to attain the most favorable outcomes.
Baseline. We conducted a comparative analysis of our proposed methodology against several contemporary baselines, which encompass regularization-based techniques such as oEWC [43], Synaptic Intelligence (SI) [65], Learning without Forgetting (LWF) [31], Classifier-Projection Regularization (CPR) [6], and Gradient Projection Memory (GPM) [42]. Additionally, we incorporated Bayesian approaches like Natural Continual Learning (NCL) [21], architecture-centric methods such as HAT [44], and memory-driven strategies including ER [10], A-GEM [9], GSS [2], DER++ [4], and HAL [8]. Moreover, we also evaluated the most recent Refresh Learning [54] paradigm within our comparative framework.
4.2 Comparison Results
All experiments were conducted with a memory capacity of 500. For each dataset, we executed the experiments 10 times and computed the mean accuracy along with the standard deviation. The overall accuracy for both Task-IL and Class-IL scenarios is summarized in Table.3.
Memory-based approaches, including ER, A-GEM, GSS, DER++, and HAL, leverage memory buffers to replay past data and mitigate catastrophic forgetting. These methods generally outperform regularization-based methods, particularly in the Task-IL scenario. For example, ER achieves 93.61% accuracy on CIFAR-10 under the Task-IL setting, while DER++ and Refresh Learning improve this the accuracy to 93.88% and 94.44%, respectively. However, in Class-IL scenarios, these methods still struggle, with ER achieving 57.74% and DER++ improving to 72.70% on CIFAR-10. While these methods perform better than regularization-based approaches, showing that the memory-based methods are suitable for addressing network forgetting in continual learning.
Our proposed MLFMM method achieves the best performance across all datasets and scenarios. For CIFAR-10, MLFMM achieves the highest accuracy of 75.29% under the Class-IL setting, significantly outperforming previous best-performing methods such as DER++ (72.70%) and Refresh Learning (73.88%). Similarly, MLFMM achieves the highest accuracy of 94.94% under the Task-IL setting, slightly surpassing Refresh Learning (94.44%). For the CIFAR-100 and Tiny-ImageNet, MLFMM consistently outperforms other methods. For example, MLFMM achieves 40.71% accuracy on CIFAR-100 under the Class-IL setting, compared to 39.10% by Refresh Learning and 36.37% by DER++. For Tiny-ImageNet, MLFMM achieves 18.38% accuracy on the Class-IL setting and 53.09% accuracy on the Task-IL setting, demonstrating its scalability to more complex datasets.
Table 3: The average accuracy of various models on 3 datasets.
@@ 4.2 Comparison Results
ER 57.74 ± 0.27 93.61 ± 0.27 20.98 ± 0.35 73.37 ± 0.43 9.99 ± 0.29 48.64 ± 0.46
DER++ 72.70 ± 1.36 93.88 ± 0.50 36.37 ± 0.85 75.64 ± 0.60 19.38 ± 1.41 51.91 ± 0.68
refresh 73.88 ± 1.16 94.44 ± 0.39 39.10 ± 0.65 76.80 ± 0.31 16.16 ± 0.72 53.36 ± 1.25
MLFMM
75.29 ± 0.75
94.94 ± 0.41
40.71 ± 0.61
76.86 ± 0.82
18.38 ± 0.31
53.69 ± 0.93
@@
Backward Transfer Analysis. Backward Transfer (BWT) measures the effect of learning new tasks. Negative values indicate a decline in performance on earlier tasks after training on new tasks, highlighting the challenge of catastrophic forgetting in continual learning. Table 4 presents the BWT scores of various methods across three datasets: CIFAR-10, CIFAR-100 and Tiny-ImageNet, in both Class-IL and Task-IL scenarios. From the results of Table 4, we can observe that the fine-tuning achieves the worst BWT across all datasets, with highly negative values in both Class-IL and Task-IL scenarios. For instance, fine-tuning achieves a BWT of $-96.39\%$ and $-46.24\%$ for Class-IL and Task-IL on CIFAR-10, respectively. This confirms that without strategies to mitigate forgetting, the network overwrites its knowledge of previous tasks while learning new ones.
Regularization-based approaches, including A-GEM and GSS slightly improve backward transfer but still struggle with catastrophic forgetting. For example, AGEM achieves $-94.01\%$ BWT in Class-IL for CIFAR-10, which is only marginally better than finetuning.
Memory-based methods such as HAL, ER, and DER++ perform better in reducing BWT. Among these, DER++ and Refresh achieve -22.38 and -22.03 BWT values on CIFAR-10 under the Class-IL setting, demonstrating notable improvements compared to other baselines. In addition, the proposed MLFMM achieves the best backward transfer across all datasets and scenarios. For example: On CIFAR-10, OW-MMD achieves a backward transfer of -20.61 in Class-IL and -3.0 in Task-IL, outperforming both DER++ and Refresh. For CIFAR-100, OW-MMD achieves -42.99 and -13.8 under the Class-IL and Task-IL setting, respectively, which are significantly better than DER++ and other baselines. For the more complex dataset such as Tiny-ImageNet, OW-MMD also demonstrates better backward transfer with-59.31 and -27.73 under the Class-IL and Task-IL settings, outperforming most baselines such as DER++ and Refresh.
@@
finetuning -96.39±0.12 -46.24±2.12 -89.68±0.96 -62.46±0.78 -78.94±0.81 -67.34±0.79
HAL -62.21±4.34 -5.41±1.10 -49.29±2.82 -13.60±1.04 - -
AGEM -94.01±1.16 -14.26±1.18 -88.5±1.56 -45.43±2.32 -78.03±0.78 -59.28±1.08
GSS -62.88±2.67 -7.73±3.99 -82.17±4.16 -33.98±1.54 - -
ER -45.35±0.07 -3.54±0.35 -74.84±1.38 -16.81±0.97 -75.24±0.76 -31.98±1.35
DER++ -22.38±4.41 -4.66±1.15 -53.89±1.85 -14.72±0.96 -64.6±0.56 -27.21±1.23
refresh -22.03±3.89 -4.37±1.25 -53.51±0.70 -14.23±0.75 -63.90±0.61 -25.05±1.05
MLFMM -20.61±0.45 -3.0 ± 0.48 -42.99±1.23 -13.8±0.84 -59.31±0.64 -27.73±1.0
@@ 4.3 Analysis Results
Forgetting Curve. Forgetting rates highlights the extent of knowledge loss over time. In this analysis, we investigate the forgetting curves for several methods.The forgetting curves for Class-IL and Task-IL settings are shown in 4, which demonstrate significant differences in performance across various methods.
For the class-incremental learning scenario, methods such as DER, DER++, and ER demonstrate relatively slow forgetting, with the forgetting rate steadily increasing as more tasks are learned, but remaining at manageable levels. Notably, DER and DER++ show smaller increments in forgetting compared to ER, indicating the effectiveness of their more advanced replay strategies combined with knowledge distillation and regularization. In contrast, oEWC and LWF exhibit a significant rise in forgetting, particularly in later tasks, highlighting the challenges of using regularization methods alone without the aid of rehearsal or distillation. The forgetting rates for these methods indicate substantial memory loss, as their performance on earlier tasks degrades significantly when new tasks are introduced.
In the task-incremental learning setup, the methods show a similar trend, although the forgetting rates are generally lower compared to the class-incremental scenario due to the more explicit task boundaries. DER and DER++ continue to perform well, showing relatively stable forgetting curves with only slight increases in forgetting over time. ER also performs reasonably well, though it demonstrates a slightly higher increase in forgetting rates as the number of tasks increases. The oEWC and LWF methods again suffer from greater forgetting, with their forgetting rates increasing more rapidly as new tasks are added. However, our MLFMM method shows the best performance across both scenarios, with its forgetting curve remaining relatively flat throughout the learning process. This indicates that MLFMM, supported by its adaptive regularization optimization, is highly effective in mitigating catastrophic forgetting, achieving stability even as new tasks are added.
In summary, the analysis of the forgetting curves reveals that while methods such as DER, DER++, and ER provide effective solutions to mitigate forgetting, oEWC and LWF struggle to maintain task performance over time. The proposed MLFMM framework, however, stands out for its ability to minimize forgetting effectively across both class-incremental and task-incremental learning scenarios. This highlights the potential of combining multi-level feature matching with adaptive regularization to tackle the challenge of catastrophic forgetting in continual learning
(a) Class-IL
(b) Task-IL
Figure 4: The forgetting curves for Class-Incremental Learning (a) and Task-Incremental Learning (b) on CIFAR10.
@@
(a) CIFAR-10
(b) TinyImageNet
Figure 5: Evolution of Weights Across Tasks for Different Layers. The figure illustrates how the weights assigned to different network layers evolve as tasks progress, showing the increasing importance of shallow layers and the decreasing trend in deep layers.
The evolution of these weights across tasks is illustrated in Fig 5. From these results, we can find several trends, summarized in the following :
@@
Figure 6: Final Weight Distribution Across Layers After Completing All Tasks. The bar chart shows the final distribution of weights for different layers of the network across datasets, highlighting that the optimal weight distribution varies between different datasets. This illustrates the different levels of importance assigned to each layer depending on the characteristics of the dataset.
@@
The bar graph in Figure 6 illustrates the final distribution of adaptive weights between the different layers of the network after completing all tasks. It is evident that the optimal weight distribution varies significantly between datasets such as CIFAR-10 and Tiny-ImageNet. For CIFAR-10, the network tends to assign higher weights to certain shallow layers, indicating their importance in generalizing across tasks. On the other hand, for Tiny-ImageNet, the weight distribution is more concentrated towards specific layers, reflecting the dataset’s complexity and the unique feature extraction requirements. These differences highlight how the network adapts its internal representations depending on the characteristics of the dataset, dynamically adjusting the importance of each layer to optimize knowledge retention and transfer in a continual learning setting.
@@
Distance Function Analysis. Table 5 compares the performance of different distance functions and the effect of using optimal weights in our experiments on CIFAR-10 with a buffer size of 500. The first column represents the distance function used, while the second column indicates whether adaptive regularization optimization term (AROT) were applied, with "F" denoting no use of adaptive weights and "T" indicating their use.
@@
From the results, it is evident that using the Maximum Mean Discrepancy (MMD) distance function yields the best performance in both the Class-IL and Task-IL scenarios. Specifically, the configuration with MMD and adaptive weights achieves the highest accuracy. Additionally, we observe that for the same distance function, the use of adaptive weights consistently outperforms the fixed weight configurations. This demonstrates the effectiveness of both MMD as a distance function and the benefits of applying adaptive weights to improve performance during continual learning.
Table 5: Comparison of distance functions and the effect of optimal weights on CIFAR-10 with buffer size 500
dist_fun AROT Class-IL Task-IL
cos F 74.62±0.79 94.53±0.18
cos T 74.89±0.86 94.67±0.27
L2 F 75.12±0.74 94.88±0.33
L2 T 75.23±0.94 94.56±0.15
mmd F 74.69±1.07 94.51±0.24
mmd T 75.29±0.75 94.94±0.41
Buffer Size.
Table 6: The average accuracy of various buffer size on CIFAR10.
Buffer Size 200 5120
Method Class-IL Task-IL Class-IL Task-IL
GEM 25.54±0:76 90.44±0.94 25.26±3.46 95.55±0.02
iCaRL 49.02±3.20 88.99±2.13 55.07±1.55 92.23±0.84
FDR 30.91±2.74 91.01±0.68 19.70±0.07 94.32±0.97
HAL 32.36±2.70 82.51±3.20 59.12±4.41 88.51±3.32
A-GEM 20.04±0.34 83.88±1.49 21.99±2.29 90.10±2.09
GSS 39.07±5.59 88.80±2.89 67.27±4.27 94.19±1.15
ER 44.79±1.86 91.19±0.94 82.47±0.52 96.98±0.17
DER++ 64.88±1.17 91.92±0.60 85.24±0.49 96.12±0.21
refresh 65.39±1.01 92.80±0.42 85.98±0.43 96.43±0.11
MLFMM 67.74±1.06 92.97±0.37 86.33±0.37 96.61±0.13
@@
Table 6 shows the performance of our proposed MLFMM method with different buffer sizes and compares it with several other continual learning methods. We evaluate the accuracy of each method on the CIFAR10 dataset under both Class-IL and Task-IL settings, using buffer sizes of 200 and 5120.
For the smaller buffer size (200), MLFMM achieves the highest accuracy in both Class-IL and Task-IL settings. Specifically, MLFMM outperforms other methods with a Class-IL accuracy of 67.74% and a Task-IL accuracy of 92.97%, which is better than any other method in this scenario. Other methods, such as GEM and A-GEM, perform relatively poorly with this smaller buffer, showing significantly lower accuracies, especially in the Class-IL setting. MLFMM’s superior performance indicates its ability to leverage smaller memory buffers effectively, which is crucial for real-world applications with limited memory.
When we increase the buffer size to 5120, MLFMM continues to perform well, although its performance advantage over other methods becomes less pronounced. In this larger buffer setting, ER and DER++ achieve slightly higher Task-IL accuracies (96.98% and 96.12%, respectively) than MLFMM’s 96.61%, while MLFMM still outperforms other methods in Class-IL. This demonstrates that MLFMM remains competitive with larger buffer sizes and continues to offer strong performance, especially in the Class-IL scenario.
In conclusion, MLFMM excels in both small and large buffer settings, particularly in Class-IL scenarios. It offers a balanced trade-off between performance and memory usage, outperforming many popular continual learning methods across different buffer sizes. These results highlight the potential of MLFMM for continual learning tasks with constrained memory budgets.
Training Time Comparison. We analyze the training times of our proposed MLFMM method in comparison with other baseline methods. The training time data for each method is shown in Figure 7. As observed, the training time for MLFMM is notably higher than most of the other methods. Specifically, MLFMM takes approximately 143 minutes, which is significantly longer than methods such as oEWC (28 minutes), LWF (32 minutes), and ER (40 minutes). However, the method still performs competitively when considering the effectiveness of the results, demonstrating a balance between computational cost and model performance. The GSS method, while highly effective, has the longest training time at 300 minutes, which suggests a trade-off between performance and efficiency. Despite the higher time cost, the MLFMM method achieves promising results in terms of accuracy and memory retention.
Figure 7: Training time comparison of different models (in minutes). The bar plot shows the runtime for each model during training, indicating significant variation in the computational cost across different methods.

### v4

@@ 4.1 Experiment Setup
Implementation Details. We employ ResNet18 as our foundational architecture, comprising five layers dedicated to feature extraction. To derive the adaptive regularization optimization term, we initialized five adaptive weights randomly, sourced from a uniform distribution. Our methodology extends the code from Refresh Learning [65], integrating our approach within their established framework. We utilized the hyperparameters specified in their implementation and conducted a grid search to ascertain the optimal value for the $\gamma$ hyperparameter, aiming to attain the most favourable outcomes.
Baseline. We conducted a comparative analysis of our proposed methodology against several contemporary baselines, which encompass regularization-based techniques such as oEWC [40], Synaptic Intelligence (SI) [24], Learning without Forgetting (LWF) [37], Classifier-Projection Regularization (CPR) [66], and Gradient Projection Memory (GPM) [67]. Additionally, we incorporated Bayesian approaches like Natural Continual Learning (NCL) [68], architecture methods such as HAT [45], and memory-driven strategies including ER [28], A-GEM [15], GSS [32], DER++ [14], and HAL [33]. Moreover, we also evaluated the most recent Refresh Learning [65] paradigm within our comparative framework.
4.2 Comparison of Results
All experiments were conducted with a memory capacity of 500. For each dataset, we executed the experiments 10 times and computed the mean accuracy along with the standard deviation. The overall accuracy for both Task-IL and Class-IL scenarios is summarized in Table 3.
Table 3: The average accuracy of various models on 3 datasets with buffer size 500.
@@ 4.2 Comparison of Results
ER 57.74±0.27 93.61±0.27 20.98±0.35 73.37±0.43 09.99±0.29 48.64±0.46
DER++ 72.70±1.36 93.88±0.50 36.37±0.85 75.64±0.60 18.90±0.09 51.84±0.47
DER++ + Refresh 73.88±1.16 94.44±0.39 39.10±0.65 76.80±0.31 16.16±0.72 53.36±1.25
OWMMD (Ours) 75.29±0.75 94.94±0.41 41.75±0.58 76.99±0.38 19.18±0.31 53.69±0.93
In comparison to fine-tuning and Joint train, OWMMD excels because these baselines are unable to effectively address catastrophic forgetting when exposed to sequential tasks. Fine-tuning shows poor performance due to its inability to retain past knowledge when exposed to new tasks, resulting in a steady decline in accuracy as more tasks are introduced. Joint train, while providing a notable improvement, still struggles in Class-IL settings, where it does not exhibit the task-specific adaptation necessary for continual learning. In contrast, OWMMD integrates not only MLFMM and AROT but also memory mechanisms, which allow the model to not only adapt to new tasks but also retain essential knowledge from previously learned tasks.
While regularization-based methods such as SI, LWF, and CPR improve performance over fine-tuning, they still fall short in preventing forgetting over a long sequence of tasks. These methods attempt to regularize the model parameters, but their capacity to effectively manage the trade-off between learning new tasks and retaining old ones is limited. For instance, while SI and CPR perform reasonably well in Task-IL settings, they still underperform in Class-IL scenarios due to their inability to fully accommodate the growing complexity of continual task learning. On the other hand, OWMMD surpasses these methods by utilizing an adaptive regularization term, ensuring that each model layer contributes appropriately to both new and old tasks, thus resulting in improved performance across all datasets, especially in Class-IL settings.
Memory-based methods such as ER and A-GEM are designed to mitigate forgetting by storing and reusing past experiences. While ER performs well in Task-IL, where it can replay past data to avoid forgetting, and A-GEM constrains updates to avoid drastic shifts in learned knowledge, both methods still face limitations in handling large-scale, complex datasets. OWMMD outperforms both ER and A-GEM, particularly in Task-IL, due to its ability to dynamically adapt the regularization weights during training. This ensures that the model does not overfit to old tasks while still retaining key knowledge when presented with new tasks.
When comparing OWMMD to more advanced methods like DER++, Refresh Learning, and HAT, our approach consistently delivers superior results, particularly on CIFAR-10 and CIFAR-100 in the Class-IL setting. While DER++ and Refresh Learning both exhibit competitive performance, OWMMD surpasses them due to its more robust handling of task-specific regularization and memory management. DER++ shows a strong performance across all datasets, but OWMMD outperforms it on CIFAR-10 and CIFAR-100, where our adaptive regularization and memory strategy provide more effective balancing of knowledge retention. Refresh Learning performs well, but it does not match the performance of OWMMD on these datasets, particularly in the Task-IL scenario, where our method achieves a higher accuracy.
In summary, OWMMD demonstrates its superiority over a wide range of baselines due to its innovative integration of adaptive regularization and memory mechanisms. These features allow our model to effectively mitigate catastrophic forgetting, retain crucial knowledge, and adapt to new tasks, outperforming both classical methods and more recent continual learning frameworks across multiple datasets and settings.
@@
Backward Transfer Analysis. Backward Transfer (BWT) measures the effect of learning new tasks. Negative values indicate a decline in performance on earlier tasks after training on new tasks, highlighting the challenge of catastrophic forgetting in continual learning. Table 4 presents the BWT scores of various methods across three datasets: CIFAR-10, CIFAR-100 and Tiny-ImageNet, in both Class-IL and Task-IL scenarios.
In the case of fine-tuning, the BWT values are substantially negative across all datasets, with CIFAR-10 and Tiny-ImageNet exhibiting values as low as -96.39 and -78.94, respectively. This suggests that fine-tuning on new tasks causes significant forgetting of previously learned tasks. Similarly, HAL and A-GEM, while better than fine-tuning, still exhibit considerable negative BWT, especially in Class-IL settings, with HAL showing -62.21 for CIFAR-10 and A-GEM showing -94.01 for CIFAR-10. These values indicate that these methods fail to effectively mitigate forgetting when new tasks are introduced, resulting in a performance drop on earlier tasks.
In contrast, OWMMD demonstrates remarkably lower BWT values, with the best performance seen in Task-IL settings, where the BWT is -3.0 on CIFAR-10. The method significantly reduces backward transfer on both CIFAR-100 and Tiny-ImageNet, with BWT values of -42.99 and -59.31, respectively, compared to higher values in other methods. This indicates that the adaptive regularization mechanism of OWMMD helps minimize the degradation of previously learned knowledge, allowing the model to retain crucial information even after the introduction of new tasks.
Other methods like ER and DER++ also show some improvement in backward transfer, but they still suffer from a noticeable decline in performance on earlier tasks. Refresh Learning demonstrates strong results, especially on CIFAR-10 and CIFAR-100, but OWMMD still outperforms it with lower BWT values, highlighting the advantages of our proposed adaptive regularization strategy in mitigating forgetting.
@@
fine-tuning -96.39±0.12 -46.24±2.12 -89.68±0.96 -62.46±0.78 -78.94±0.81 -67.34±0.79
HAL -62.21±4.34 -05.41±1.10 -49.29±2.82 -13.60±1.04 - -
A-GEM -94.01±1.16 -14.26±1.18 -88.50±1.56 -45.43±2.32 -78.03±0.78 -59.28±1.08
GSS -62.88±2.67 -07.73±3.99 -82.17±4.16 -33.98±1.54 - -
ER -45.35±0.07 -03.54±0.35 -74.84±1.38 -16.81±0.97 -75.24±0.76 -31.98±1.35
DER++ -22.38±4.41 -04.66±1.15 -53.89±1.85 -14.72±0.96 -60.71±0.69 -29.01±0.46
DER++ + Refresh -22.03±3.89 -04.37±1.25 -53.51±0.70 -14.23±0.75 -65.07±0.73 -27.36±1.49
OWMMD(ours) -20.61±0.45 -03.00±0.48 -42.99±1.23 -13.80±0.84 -59.31±0.64 -27.73±1.00
Overall, the results emphasize that OWMMD stands out as the most effective approach in terms of minimizing backward transfer, addressing the challenge of catastrophic forgetting, and ensuring that the model can learn new tasks without significant degradation in the performance of previously learned tasks.
4.3 Comparison of Results for Complex Datasets
In the comparison on complex datasets, as shown in Table 5, OWMMD demonstrates superior performance across all three datasets (Cars-196, Cub-200, and ImageNet-R), significantly outperforming other methods. On the Cars-196 and Cub-200 datasets, both in Class-IL and Task-IL settings, OWMMD leads by a large margin, particularly in the Task-IL setting, with accuracies of 52.53% and 48.69% respectively, far surpassing other methods like DER++ (36.59% and 37.68%) and ER (26.06% and 28.71%). On the ImageNet-R dataset, OWMMD similarly shows strong robustness, with accuracies of 8.93% in Class-IL and 30.39% in Task-IL, outperforming DER++ (5.82% and 22.73%) and DER (3.88% and 17.17%) among others.
Table 5: Average accuracy of various models on 3 complex datasets with buffer size 500.
Method Cars-196 Cub-200 ImageNet-R
Class-IL Task-IL Class-IL Task-IL Class-IL Task-IL
LWF 07.27±0.13 11.99±0.05 04.89±0.10 09.95±0.49 04.93±0.19 11.55±0.39
oEWC 05.22±0.62 10.53±0.38 04.34±0.12 09.83±0.45 04.28±0.52 12.98±0.82
ER 05.63±0.09 26.06±1.19 07.05±0.34 28.71±0.63 03.94±0.82 19.68±3.97
DER 07.92±0.70 31.50±0.34 06.79±0.37 29.87±0.86 03.88±0.76 17.17±3.31
DER++ 09.34±0.33 36.59±1.19 13.50±3.11 37.68±3.89 05.82±0.64 22.73±2.23
DER++ + Refresh 08.70±0.57 35.63±1.00 13.59±0.74 37.79±0.66 05.47±0.49 22.50±1.12
OWMMD (Ours) 18.09±0.86 52.53±1.24 22.84±2.49 48.69±1.37 08.93±1.14 30.39±2.33
These results indicate that OWMMD not only achieves excellent results on standard datasets but also excels when faced with more challenging and complex tasks. Especially in preventing catastrophic forgetting and maintaining high accuracy, OWMMD stands out. Compared to other methods, OWMMD is particularly effective in multi-task settings, adapting better to the variability and complexity of complex datasets, offering a more efficient and robust solution for continual learning tasks.
@@ 4.4 Analysis Results
Forgetting Curve. Forgetting rates highlights the performance and knolwedge losses over time. In this experiment, we investigate the forgetting curves for several methods on the CIFAR-10, CIFAR-100 and Tiny-ImageNet datasets under both class-incremental learning (Class-IL) and task-incremental learning (Task-IL), and the results are presented in Figure 3. The empirical results demonstrate that different methods show different forgetting behaviors.
Figure 3: Forgetting curve analysis of various models on CIFAR-10 , CIFAR-100 and Tiny-ImageNet datasets under class-incremental and task-incremental learning settings.
In CIFAR-10, OWMMD outperforms all other methods under both Class-IL and Task-IL settings, showing a notably lower forgetting rate. This is particularly evident in the task progression where OWMMD maintains stable performance over time, while other methods, such as DER and ER, experience more significant forgetting. A similar trend holds for CIFAR-100 and Tiny-ImageNet, where OWMMD exhibits superior stability across tasks, demonstrating its effectiveness in retaining knowledge from earlier tasks. Furthermore, the proposed OWMMD achieves the lowest forgetting rate than other baselines at each task learning, demonstrating its strong ability to fight forgetfulness.
@@
The evolution of these weights across tasks is illustrated in Figure 4. From these results, we can find several trends, summarized in the following :
@@
Figure 4: Dynamic Adjustment Process of Layer-wise Adaptive Weights Across Tasks.
Figure 5: The distribution of the final adaptive weights after learning all tasks.
@@
Figure 5 illustrates the final distribution of adaptive weights between the different layers of the network after completing all tasks. It is evident that the optimal weight distribution varies significantly between datasets such as CIFAR-10 and Tiny-ImageNet. For CIFAR-10, the network tends to assign higher weights to certain shallow layers, indicating their importance in generalizing across tasks. On the other hand, for Tiny-ImageNet, the weight distribution is more concentrated towards specific layers, reflecting the dataset’s complexity and the unique feature extraction requirements. These differences highlight how the network adapts its internal representations depending on the characteristics of the dataset, dynamically adjusting the importance of each layer to optimize knowledge retention and transfer in a continual learning setting.
@@
Distance Function Analysis. Table 6 compares the performance of different distance functions and the effect of using optimal weights in our experiments on CIFAR-10 with a buffer size of 500. The first column represents the distance function used, while the second column indicates whether the adaptive regularization optimization term (AROT) was applied, with "False" denoting no use of adaptive weights and "True" indicating their use.
@@
From the results, it is evident that using the Maximum Mean Discrepancy (MMD) distance function yields the best performance in both the Class-IL and Task-IL scenarios. Specifically, the configuration with MMD and adaptive weights achieves the highest accuracy. Additionally, we observe that the use of adaptive weights consistently outperforms the fixed weight configurations. This demonstrates the effectiveness of both MMD as a distance function and the benefits of applying adaptive weights to improve performance during continual learning.
Table 6: Comparison of different distance functions and the effect of adaptive regularization optimization term (AROT) on CIFAR-10 with buffer size 500. The first column shows the distance used (cos, L2, or MMD). The second column indicates whether the adaptive weights (AROT) were applied, where "True" denotes the use of adaptive weights and "False" means no adaptive weights were applied. The last two columns display the performance results for Class-IL and Task-IL scenarios, including mean accuracy with standard deviation. The highest performance in each case is marked in bold.
Distance AROT Class-IL Task-IL
Cosine False 74.62±0.79 94.53±0.18
Cosine True 74.89±0.86 94.67±0.27
L2 False 75.12±0.74 94.88±0.33
L2 True 75.23±0.94 94.56±0.15
MMD False 74.69±1.07 94.51±0.24
MMD True 75.29±0.75 94.94±0.41
@@
Buffer Size. Table 7 shows the performance of our proposed OWMMD method with different buffer sizes compared with continual learning baselines. Evaluations on CIFAR-10 under Class-IL and Task-IL settings reveal three key observations:
With the smaller buffer size (200), OWMMD achieves the highest accuracy in both scenarios (67.74% Class-IL, 92.97% Task-IL), outperforming all compared methods. Notably, its Class-IL accuracy surpasses the second-best DER$+$Refresh by 2.35%, while maintaining a 0.17% Task-IL advantage. This demonstrates superior capability in memory-constrained environments where methods like GEM (25.54% Class-IL) and A-GEM (20.04% Class-IL) show significant degradation.
When using the 5120 buffer, OWMMD maintains competitiveness with 86.33% Class-IL accuracy (best overall) and 96.61% Task-IL accuracy. Although ER achieves marginally higher Task-IL performance (96.98%), it suffers a 14.51% accuracy drop in Class-IL compared to its Task-IL results, while OWMMD shows balanced performance with only 10.3% difference.
The progressive improvement from 200 to 5120 buffers confirms OWMMD’s effective memory utilization: Class-IL accuracy increases by 18.59% absolute (67.74% $\rightarrow$ 86.33%), significantly exceeding the average 15.2% gain of other methods. This demonstrates our method’s unique advantage in scenarios requiring flexible memory scaling without performance saturation.
Table 7: Performance comparison with buffer size 500 as baseline. Arrows (↑/↓) indicate performance changes relative to buffer size 500. Our method shows progressive improvement with larger buffers.
Method 200 500 5120
Class-IL Task-IL Class-IL Task-IL Class-IL Task-IL
GEM 25.54±0.76↓ 90.44±0.94↓ 26.20±1.26 92.16±0.69 25.26±3.46↓ 95.55±0.02↑
iCaRL 49.02±3.20↑ 88.99±2.13↑ 47.55±3.95 88.22±2.62 55.07±1.55↑ 92.23±0.84↑
FDR 30.91±2.74↑ 91.01±0.68↓ 28.71±3.23 93.29±0.59 19.70±0.07↓ 94.32±0.97↑
HAL 32.36±2.70↓ 82.51±3.20↓ 41.79±4.46 84.54±2.36 59.12±4.41↑ 88.51±3.32↑
A-GEM 20.04±0.34↓ 83.88±1.49↓ 22.67±0.57 89.48±1.45 21.99±2.29↓ 90.10±2.09↑
GSS 39.07±5.59↓ 88.80±2.89↓ 49.73±4.78 91.02±1.57 67.27±4.27↑ 94.19±1.15↑
ER 44.79±1.86↓ 91.19±0.94↓ 57.74±0.27 93.61±0.27 82.47±0.52↑ 96.98±0.17↑
DER++ 64.88±1.17↓ 91.92±0.60↓ 72.70±1.36 93.88±0.50 85.24±0.49↑ 96.12±0.21↑
DER$+$Refresh 65.39±1.01↓ 92.80±0.42↓ 73.88±1.16 94.44±0.39 85.98±0.43↑ 96.43±0.11↑
OWMMD (Ours) 67.74±1.06↓ 92.97±0.37↓ 75.29±0.75 94.94±0.41 86.33±0.37↑ 96.61±0.13↑
Training Time Comparison. In terms of training time, OWMMD demonstrates significant efficiency improvements on CIFAR-10 and CIFAR-100 compared to many baselines. For CIFAR-10, OWMMD completes training in 143 minutes—2.4× faster than GSS (341 minutes) and nearly twice as fast as GEM (280 minutes). While lightweight methods like LWF (32 min) and oEWC (28 min) require less time, OWMMD maintains a practical training duration that remains competitive with most approaches, particularly in scenarios involving multiple sequential tasks. This efficiency advantage persists on CIFAR-100, where OWMMD (149 min) trains 2.7× faster than GSS (395 min) and 4.7× faster than GEM (705 min), while retaining strong performance.
Figure 6: Training time comparison of various models on CIFAR-10/CIFAR-100 (50 epochs) and Tiny-ImageNet (100 epochs) Datasets with RTX 4090D GPU.
For Tiny-ImageNet, OWMMD requires 22.3 hours, which exceeds DER++ (6.63 hours) and LWF (2.54 hours) by 3.4× and 8.8× respectively. However, this increased computational cost is offset by OWMMD’s superior accuracy and stability in continual learning scenarios. The method’s ability to mitigate catastrophic forgetting while maintaining competitive runtime efficiency, despite the dataset’s complexity, positions it as a robust solution for continual learning tasks.